Solve by substitution method x-2y=-2...
X^2+y^2-2x+3y=4

Answer by Boreal(15235)

x-2y=-2
x=2y-2
(2y-2)^2+y^2-2(2y-2)+3y=4
4y^2-8y+4+y^2-4y+4+3y=4
5y^2-9y+4=0
(5y-4)(y-1)=0
y=4/5, 1
x=-0.4, 0
(0, 1) and (-0.4, 0.8)
check
0.16+0.64+0.8+2.4=4
0+1-0+3=4